16) Christine bought a handbag for $53.40 on sale. It was marked down 40% off the original price. How much did Christine save?
17) Joe got a promotion at work and a 15% raise. If Joe’s new salary is $48,000 a year, how much did Joe make before his raise?
18) The length of a rectangle is 3 inches more than the its width. If the perimeter of the rectangle is 30 inches, find the length and width of the rectangle.
19) The sum of the angles of a triangle is 180o. The first angle of the triangle measures double the second angle. The third angle of the triangle measures 20 degrees more than the second. Find the measure of each angle.
20) Brenda and Gillian live 1 mile apart along the same street. Brenda walks at a rate of 3 miles per hour and Gillian walks at a rate of 2.5 miles per hour. If they each leave their own house at the same time, how long, to the nearest minute, will it take for them to meet? (HINT: Remember that Distance = Rate x Time)


Found 2 solutions by rfer, mathmandotcom:
Answer by rfer(16322)

16)
53.40/.6=89
89-53.40=35.60 saved
17)
48000/1.15=41,739.13
18)
2x+2(x+3)=30
19)
2x+x+x+20=180
20)
toatl speed is 5.5 mph
1/5.5=0.183 hrs=10.98 min=11 min

Answer by mathmandotcom(31)

16) christine paid 60% of retail price or Rp
.6x=53.40 solve
x=53.40/.6
x=534/6=89
original price=$89
she paid 53.40 subtract to get savings
89.00
53.40
_----
35.60
Christine saved $35.60
17) 1.15x=48000 x+.15x=48000 where x=oringinal
x=48000/1.15
x=41,739.13
18) length=3+width
l=3+w
P=2l+2w=30
P=2(3+w)+2w sub
P=6+2w+2w
P=6+4w=30
subtracting 6 we get
4w=24 divide by 4
w=6 then l=w+3=9
w=6,l=9 answer
19) second angle=x
first angle=2x
third angle=x+20 so x+2x+x+20=180 4x=160 divide by 4 second angle=40
first=2*40=80
second=40
third=20+40=60
20) d1=3t
d2=2.5t
D=1 mile
D=d1+d2=3t+2.5t=1hr and t is in hours
5.5t=1 so t=1/5.5=.181818hrs multiply by 60 to convert to minutes
10.9minutes nearest minute 11 answer 11 minutes
I hope this helps, good luck and check your answers:)
first=